A Complete Bibliography of ACM Transactions on Asian Language Information Processing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.50

Title word cross-reference

F [120]. k [52]. N [45].

-grams [45]. -nearest [52]. -term-based [120].

2005 [88]. 2008 [155].

4 [61, 64, 59].

6 [117].

A* [67]. Abbreviations [204]. ABSPs [125]. access [65, 32, 72]. accessing [19]. Acoustic [167]. Acquiring [119, 76]. Acquisition [122]. adaptation [90, 44, 18]. Adapting [92, 120]. Adaptive [30, 167, 190, 52]. Advances [127, 72]. against [101]. Algorithm [184, 169, 14]. Algorithm-Based [169].

Aligning [84]. Alignment [208, 132, 85]. Allograph [228]. Analyses [170]. Analysis [185, 49, 145, 66, 36, 44, 227]. Anaphora [165, 75, 111]. Anaphoric [165]. anaphoricity [75]. anchor [9]. annotated [119]. Annotation [219]. Answer [181, 68]. Answering [119, 125, 66, 65, 67, 68, 102, 29]. antecedent [75]. Application [121, 210, 77, 106]. Applications [170, 6]. Approach [169, 208, 112, 46, 2, 50, 10, 229]. Approaches [181]. Arabic [165, 142, 162, 141, 140, 161, 225, 164, 144, 163, 139, 217, 220, 183, 143]. Argument [197]. Articles [202, 53, 69]. Articulation [167]. Articulation-Disordered [167]. Asian [131, 226, 62, 127]. aspects [55]. Assisted [214]. association [99]. Attention [203]. Attention-Feedback [203]. augmented [48]. AUTHOR [78]. Auto [136]. Auto-Parsed [136]. Automatic [165, 41, 144, 11, 213, 172, 105, 229]. Automatically [119, 53].

Based [149, 142, 184, 114, 169, 141, 218, 189, 209, 210, 191, 233, 168, 175, 197, 211, 203, 188, 166, 220, 183, 174, 150, 46, 73, 110, 48, 50, 36, 4, 26, 11, 15, 10, 120, 45, 71, 16, 77, 83, 47, 89, 79, 108, 13, 103, 104, 178]. Bayesian [208, 190]. Behaviors [180]. Bengali [156, 157]. Best [213, 88]. between [192, 101, 93, 227, 89, 175]. bigrams [15]. Bilingual [187, 84, 110, 85, 59]. Bilingually [133]. biological [82]. biomedical [81, 80, 79]. biomedicine [78]. Blind [157, 229]. Boolean [70]. Boosting [125]. Boundary [166]. Boundary-Based [166]. Brahmic [205]. break [11]. break-index [11]. Broadcast [147]. Building [6].

C*ST*RD [32]. Calligraphic [234, 107]. Cantonese [5, 49]. Case [178, 172]. Categories [164]. Categorization [220, 52, 50, 3]. Causal [119, 76]. Cebuano [31, 28]. challenge [65]. Challenges [140]. Character [200, 188, 150, 15, 17, 13, 107]. Character-Based [188, 150, 15]. Characters [215, 170, 222, 228]. checker [36]. Checking [198]. Chief [207]. Chinese [184, 6, 114, 136, 171, 215, 92, 177, 48, 2, 151, 200, 201, 14, 61, 18, 125, 15, 10, 179, 173, 124, 129, 71, 170, 12, 17, 204, 135, 181, 188, 178, 58, 174, 190, 150, 118, 172, 107, 234]. Chinese-English [178]. Choose [213]. chronological [69]. Chunks [209]. claim [63]. Class [164]. Classification [177, 120]. Classifier [169, 113]. CLIR [25]. Closed [164]. Closed-Class [164]. Cloud [234]. Cluster [145]. Clustering [138, 228, 110]. Co [198]. Co-Mention [198]. Collaborative [180]. Combination [142, 132, 45, 70]. Combining [206, 15]. Commentary [221, 226]. Comparable [179, 110]. Comparative [156, 129, 62]. Comparison [187, 8, 108, 12, 30]. complementary [77]. Complex [210]. Compositional [158]. compression [103]. Computational [161]. Computer [214, 51]. Computer-Assisted [214]. Concatenation [209]. conditional [34]. confix [112]. confix-stripping [112]. Conjugation [103]. Conjugation-based [103]. connective [76]. connectivity [4]. Considerations [226]. considering [63]. Constituent [175]. constrained [228]. Construction [177, 14]. Context [201, 45]. context-free [45]. contextual [15]. continuous [5]. contours [49]. Contradiction [198]. contrastive [81]. Conversion [200, 172, 4]. Convolution [126]. Corpora [119, 179, 84, 110, 85]. Corpus [192, 168, 122, 11, 56, 74]. Corpus-Based [168, 11]. Correction [56, 163]. Corrective [115]. correspondence [89]. Count [221]. Counts [221]. Coverage [197]. creation [22]. Cross [33, 192, 121, 219, 32, 16, 27, 118, 143, 96, 44, 61, 29, 58]. Cross-Language [192, 121, 143, 33, 16, 96]. Cross-Lingual [219, 32, 27, 44, 61, 29, 58]. Cultural [226]. customization [35].

Data [136, 206, 100, 101, 26]. databases [107]. days [25]. Deception [226]. Decoding [172]. Decomposition [184, 209, 17]. Decomposition-Based [184]. Deep [181]. Dependency [136, 175]. Dependent [152]. Design [205]. Detect [232]. Detection [165, 142, 180, 226, 166, 143, 21]. determination [75, 13]. Devanagari [228]. Developing [109, 174]. development [34, 22]. Diagnosis [184]. Dialectal [163]. dialogue [37]. dialogues [65]. Dictionary [163, 14]. different [227, 108]. dimensional [107]. Directed [210]. Disambiguation [149, 206, 18, 97]. Discourse [225]. discovering [10, 99]. discovery [83]. Discriminative [141, 233, 134, 188, 46, 3, 99]. Disordered [167]. distance [73]. Distortion [218]. Diverse [199]. Document [128, 129, 186, 127, 92, 16, 12, 68, 70]. document-length [92]. Documents [121, 222, 46]. Domain [124, 55, 65]. Domain-specific [55]. DUSTer [23].

edit [73]. edit-distance [73]. Editor [207]. Editor-in [207]. Editorial [221, 226, 207]. Effect [93]. effective [229, 59]. Effectiveness [222]. Efficient [211, 234, 73, 19]. Elementary [225]. embedded [10]. Emotion [87]. empirical [90, 65]. Employing [175]. engine [68]. Engineering [191]. English [232, 6, 92, 162, 177, 192, 216, 189, 61, 26, 32, 157, 179, 29, 211, 204, 178, 58, 118]. English-to-Japanese [189]. enhanced [108]. Ensemble [169]. Entailment [194, 195, 197, 199]. Entities [175, 85]. Entity [169, 209, 179, 183, 174, 24, 34, 79, 113]. Entrance [194]. entropy [48, 106]. Entry [224]. Environment [234]. equivalence [24]. ergodic [57]. Error [184, 115, 36]. Error-Corrective [115]. error-pattern [36]. Errors [162, 56]. Estimating [74]. Estimation [145]. Evaluating [194]. Evaluation [205, 155, 102, 54]. Event [116, 82]. Examinations [194]. Example [73, 47, 109]. Example-based [73, 47]. Exercise [155]. exercises [20]. Expansion [206, 229, 58]. experiment [22]. Experiments [26, 92]. Exploiting [215, 164, 197]. Expression [119]. Expressions [210, 37, 86]. Extended [123, 88]. Extension [187]. extracted [53, 69]. Extracting [24, 81, 40]. Extraction [187, 219, 160, 175, 174, 82, 105, 3, 35].

F0 [49]. factoid [102]. FAQ [55]. Fast [168]. Feature [191, 34]. Features [166, 111]. Feedback [157, 203, 229, 60]. fields [34]. filtering [54]. FIRE [155]. Flexible [60]. focused [68]. followed [75]. Framework [202, 186, 174, 150, 39, 97]. free [45]. Fully [190]. Functional [210]. fusion [16].

generalized [30]. generate [50]. Generation [130, 151, 116, 33]. Generative [188]. Genetic [169]. Global [233, 134]. GLR [7]. gram [46, 13]. grammar [36]. grammar-checker [36]. Grammars [178, 45, 7]. grams [45]. granularity [98]. grapheme [4]. graphemes [89]. Greetings [207]. Guest [221, 226].

Handle [176]. Handwriting [182]. Handwritten [203, 228, 17]. Hausdorff [30]. headline [33]. Hebrew [103]. Hidden [138, 47]. Hierarchies [121]. high [66, 107]. high-dimensional [107]. high-performance [66]. Hindi [21, 156, 33, 23, 31, 25, 26, 32, 157, 34, 30, 28, 29, 224, 27]. Hindi-English [29]. Hindi-to-English [26]. History [152]. HMM [46, 57, 16, 190]. HMM-based [16]. HMM/N [46]. HMM/N-gram-based [46]. homophone [18]. honorific [86]. HPSG [189]. HPSG-Based [189]. Hybrid [126, 118, 45]. Hypotheses [130]. Hypothesis [132].

Identification [137, 147, 170, 75]. IJCNLP [88]. image [30]. Images [234]. important [99]. Improve [119]. Improved [177, 132]. improvement [67]. Improving [165, 115, 36, 69, 99, 108]. Incorporating [185]. Incorrect [170]. Incremental [229]. independent [55]. index [11, 107]. Indexing [156, 157, 12]. Indian [154, 182, 230]. Indic [122]. Indonesian [112]. Inducing [148]. Induction [190, 34]. Inductive [184]. Inference [196]. Inferential [95]. Information [223, 154, 160, 175, 204, 118, 143, 96, 14, 65, 4, 53, 81, 5, 32, 71, 77, 38, 93, 35, 95, 59, 72, 94, 58]. Information-Extraction [160]. Informative [116]. Informed [172]. Inspired [151]. Integrating [121, 188]. Inter [166]. Inter-Syllable [166]. Interactive [107, 31]. Interruption [166]. Introduction [171, 131, 88, 43, 154, 117, 161, 38, 193, 51, 72, 139, 94, 127]. IR [157, 70]. Iraqi [162]. Isolated [203]. Issue [171, 131, 154, 161, 193, 139, 127, 43, 38, 51, 72, 94, 117]. issues [64]. Italian [217]. ITG [216, 227].

Japanese [215, 192, 216, 66, 189, 210, 93, 67, 195, 72, 86, 211, 135]. Japanese-to-English [211]. Jawi [222]. Journal [153, 231].

K-ToBI [11]. Kernel [126, 175]. Kernel-Based [175]. Keyboards [205]. Knowledge [185, 197, 148, 76, 85, 83, 79]. Korean [232, 41, 36, 37, 61, 191]. Kurdish [223].

Labeled [206]. Labeling [126, 218, 173, 138, 91]. labels [87]. Language [130, 171, 131, 192, 140, 121, 154, 173, 160, 152, 213, 139, 122, 135, 127, 178, 118, 143, 100, 33, 2, 43, 90, 96, 39, 44, 18, 45, 16, 35, 7, 227, 95, 20, 108, 94, 13, 104]. Languages [159, 156, 192, 213, 226, 217, 51, 230, 62]. large [107]. Latent [192, 44]. learned [59]. Learners [232]. Learning [202, 184, 206, 195, 204, 181, 48, 111, 82, 8, 56, 99]. Left [201]. Legal [202]. Length [145, 92]. Lessons [59]. Leveraging [199]. Lexical [44, 145, 197, 135, 134, 199, 79]. lexicographic [109]. Lexicon [187]. Light [148]. Lightweight [125]. limited [24]. Linear [233]. Lingual [219, 44, 61, 32, 29, 27, 58]. Linguistic [191, 227, 22]. Linguistically [151, 210]. Linguistically-Directed [210]. Linguistics [161]. literature [81, 80]. Local [233, 204]. Log [233]. Log-Linear [233]. Logic [196]. Logical [202]. Lookup [163]. Low [213]. Low-Resource [213].

Machine [131, 215, 162, 141, 216, 218, 132, 176, 158, 133, 195, 135, 134, 178, 73, 8, 56, 89]. machine-learning [8, 56]. Machines [120]. Making [31]. Malay [222]. management [78]. Mandarin [46]. Manuscript [234]. Marathi [156, 157]. marker [76]. Markov [47, 196, 106]. maximum [48, 106]. maximum-entropy [48]. ME [79]. ME-based [79]. meaning [40]. Meaningful [3]. Measuring [101]. Mechanism [184]. Mention [142, 198, 143]. metadata [50]. Method [232, 206, 209, 63, 56, 13]. methodology [3]. Methods [125, 233, 8]. Minimal [201]. Mining [208, 146, 179, 80, 100, 9, 83, 78]. MIRACLEs [31]. miserly [26]. Mispronunciations [232]. mixture [87]. Mobile [234]. modality [56]. Model [114, 218, 151, 191, 152, 90, 44, 105, 18, 45, 16, 47, 89, 108, 106]. Model-Based [191]. Modeling [204, 181, 178, 228, 100, 2, 43, 10, 49]. Models [128, 141, 129, 188, 167, 96, 85, 12, 95, 108, 62, 87, 70]. Monolingual [132, 61, 59, 62]. month [21]. Morpheme [4, 108]. Morpheme-based [4, 108]. Morphemes [148]. Morphology [142, 122]. Morphology-Based [142]. morphophonemic [4]. Motivated [133]. MRD [149]. MRD-Based [149]. MT [26]. Multi [68, 186, 16]. Multi-answer-focused [68]. Multi-Document [186, 68]. multi-scale [16]. Multidocument [116]. multilingual [62]. multimodal [37]. multiple [85, 7, 69].

N [13]. N-gram [13]. Name [114, 124, 147, 28]. Named [169, 209, 179, 175, 113, 183, 24, 85, 34, 79]. Natural [140, 139, 39, 7, 94]. nearest [52]. negation [81]. neighbor [52]. network [57]. Networks [196]. neural [57]. News [147, 41, 53]. newspaper [69]. noise [101]. Nominal [173]. Non [165, 204, 227]. Non-Anaphoric [165]. non-ITG [227]. Non-Local [204]. nonstationary [106]. normalization [92]. Normalizing [210]. Noun [152]. nouns [40]. novel [65]. NTCIR [117, 61, 64, 59]. NTCIR-4 [61, 64, 59]. NTCIR-6 [117].

OCR [30]. Offline [182, 17]. Online [198, 203, 228]. ontology [42]. OOV [176, 58]. open [65]. open-domain [65]. order [227]. Ordering [216, 211, 69]. Organization [114]. oriental [51]. Ought [221]. Overlapping [201].

Pairs [179, 181, 105, 227]. papers [88]. Paragraphs [202]. parallel [85]. Paraphrasing [210]. Parsed [136]. parser [48]. Parsing [136, 216, 36, 7]. Part [154, 164, 190, 8]. Part-of-Speech [164, 190, 8]. partial [36]. patent [63]. Patents [120]. Pattern [198, 111, 36]. Patterns [119, 167, 4, 81]. PCFG [104]. N-gram-based [46]. performance [66, 108]. Persian [57]. Personalized [167, 234]. Perspective [180]. Perspectives [221, 226, 212]. Phase [202, 82]. phoneme [4]. phonemes [89]. phonetic [4, 105, 19]. Phonologically [170]. Phrase [141]. Phrase-Based [141]. phrases [91]. Phrasing [191]. Pivot [187, 213, 135]. pivoted [92]. Platform [200]. Point [166]. Poor [160, 61, 230]. porting [23]. pos [106]. pos-tagging [106]. Position [174]. Post [216, 211, 58]. Post-Ordering [216, 211]. post-translation [58]. postprocessing [15]. Pragmatic [226]. Predicate [197]. Predicate-Argument [197]. Predicates [210, 173]. Prediction [232, 224, 11]. Preface [64]. prepositional [91]. prepositions [40]. Preprocessing [189]. Prior [185]. Probabilistic [129, 234, 70]. Probability [145]. Problem [176]. Processing [171, 140, 139, 38, 51, 72, 94]. Professional [200]. Projection [219]. Prologue [1]. Pronouns [165]. Pronunciation [232]. Propagation [143]. Proposal [63, 65]. Prosodic [191, 166]. pseudo [60]. pseudo-relevance [60]. psychological [86]. Punctuation [151].

QAT [125]. queries [9, 7]. query [92, 96, 97, 229, 58]. Question [119, 125, 181, 66, 65, 67, 68, 102, 29]. question-answering [67, 68, 29]. Questions [119]. Qur’an [145].

radical [17]. random [34]. Ranking [129]. Rapid [23, 34, 35]. reasoning [98, 94]. Recognition [123, 169, 209, 115, 194, 182, 197, 199, 167, 183, 5, 15, 34, 79, 108, 17, 113, 87]. Recognize [195]. Recognizing [196]. Reduction [191, 47]. Reexamination [149]. referring [37]. Regional [182]. Related [180, 80]. Relation [119, 219, 175, 174]. Relation-annotated [119]. Relational [192]. Relations [136, 10]. relationships [93]. Relevance [157, 181, 60]. Relevant [121, 222, 71]. reliability [102]. RENAR [183]. Reordering [177, 227]. representation [11]. Reranking [115]. research [22]. Resolution [165, 37, 124, 39, 75, 111]. Resource [159, 160, 213, 230, 22]. Resource-Poor [160, 230]. Resource-Scarce [159]. Resources [199, 24, 61]. restriction [74]. results [61]. Retrieval [128, 223, 154, 180, 127, 118, 234, 46, 73, 96, 14, 61, 71, 16, 12, 77, 63, 93, 95, 59, 55, 27, 70, 58]. Retrieving [121, 222]. Rich [61, 111]. Right [201]. RITE [193]. Robust [203]. Role [126, 173, 91]. Roman [137]. rough [100]. Rule [210, 183]. Rule-Based [210, 183]. sampling [60]. satisfactoriness [74]. scale [16]. Scarce [159]. scenario [26]. SCO [125]. SCO-QAT [125]. Script [137, 15]. Scripts [205, 182]. Search [156, 192, 138, 31, 67, 62]. section [88]. Segmentation [142, 201, 133, 203, 188, 150]. Segments [123]. Selection [134, 3]. selectional [74]. selective [60]. Semantic [126, 173, 175, 181, 91, 44, 87, 42]. semantically [80]. Semi [206]. Semi-Supervised [206]. semisupervised [110]. Sense [149, 206]. senses [84]. sentence [73, 47]. Sentences [184, 10, 69]. Sentiment [185]. separable [87]. Separation [164]. Sequence [218]. sequential [99]. set [100, 13]. Short [136]. sign [104]. Similar [170]. similarity [101, 105]. size [92]. smooth [57]. SMT [177, 233]. Soft [205]. Solutions [140]. sorting [109]. Soundbite [147]. sources [85]. spam [54]. Spanish [217]. Speaker [191, 147, 167]. Speaker-Adaptive [167]. Special [171, 131, 117, 161, 193, 139, 127, 43, 38, 51, 64, 72, 94, 88, 154]. specific [55]. Speech [115, 164, 144, 147, 167, 166, 190, 57, 5, 49, 8, 108]. Speech-to-Text [144]. Spelling [163]. Splitting [225]. Spoken [130, 128, 129, 127, 46, 16]. Spontaneous [166]. Stacking [191]. Stage [130, 63]. Standard [186]. State [153, 231]. Statistical [187, 96, 216, 218, 151, 133, 2, 43, 85, 97, 104, 54]. statistics [99, 13]. Stemmer [168, 222]. Stemming [112, 230]. stochastic [18, 45]. Strategies [156, 12, 59]. strategy [52]. stripping [112]. stroke [228]. Structure [114, 197, 174, 77, 63]. Structure-Based [114, 197, 174]. Structures [202]. Study [156, 129, 178, 172, 90, 65, 62, 86, 106]. Sub [157]. Sub-Word [157]. Sublexical [176]. Substitution [178]. Summarization [129, 186, 68]. Summary [116]. Supervised [206, 185, 219]. Support [120]. Sura [145]. Surprise [28, 35, 20]. Survey [182]. Syllable [201, 166]. Synonymous [146]. syntactic [111]. Syntax [211, 178]. Syntax-Based [211, 178]. synthesizer [57]. System [132, 214, 160, 224, 197, 183, 66, 37, 26, 35, 67, 108, 29].

Tagging [164, 217, 150, 41, 8, 106]. Taiwanese [19, 104]. TALIP [221, 226, 212]. tame} [76]. Tamil [108, 203]. Task [18, 106]. Technique [118, 19]. Techniques [182, 100, 54]. technologies [65]. technology [22]. Temporal [217, 53, 10, 38, 40]. term [3, 120]. Terminology [83]. Terminology-based [83]. Terms [204, 71, 80]. Testing [184]. Text [144, 122, 224, 220, 52, 57, 76, 3, 49, 9, 87, 78]. text-to-speech [49]. Texts [225, 195, 196, 198, 50, 61, 103]. Textual [194, 195, 197, 199]. Thai [137, 8]. thematic [50]. theory [100]. thesaurus [74]. thirty [25]. three [8]. Time [217, 39, 42, 98]. TIMEX2 [41]. ToBI [11]. Tokenization [164]. tone [11, 5, 123]. Topic [128, 110, 98, 77, 152, 21, 53]. Topic-Dependent [152]. Topic-structure-based [77]. Topics [138]. tracking [21, 110, 53, 98]. Training [233]. Transcription [128, 144]. transcriptions [19]. Transcripts [147]. Transfer [104, 26]. Transfer-based [104, 26]. transformation [48]. transformation-based [48]. Translation [114, 130, 131, 215, 162, 141, 216, 218, 132, 176, 214, 189, 9, 133, 213, 195, 211, 135, 134, 178, 73, 96, 97, 56, 104, 58]. Translations [176]. translingual [6, 31, 24]. Transliteration [159, 208, 158, 105, 89]. Transliterations [146, 101]. Tree [126, 175, 178]. Treebank [172]. Treebanks [172]. triggers [44]. TRUES [123]. Two [202, 130, 82, 125, 63]. Two-Phase [202, 82]. Two-Stage [130, 63]. typologies [227].

Understanding [115]. Unified [173, 186, 150, 2]. Units [225]. University [194]. unknown [3]. unknown-word [3]. Unsupervised [190]. Update [186]. Urdu [109, 160]. use [62]. useful [65]. Usefulness [53]. User [180]. Using [184, 126, 100, 123, 136, 192, 50, 176, 116, 5, 191, 147, 198, 86, 148, 134, 167, 166, 178, 228, 84, 73, 57, 76, 115, 4, 61, 11, 85, 34, 16, 9, 30, 67, 68, 47, 79, 113, 55, 87, 104]. Utilization [195].

Validating [198]. Variability [191]. Variant [124]. Variants [232]. Vector [120]. verb [56]. Verbal [173]. verification [82]. via [3, 60, 172]. Vietnamese [217, 113]. Visually [170]. Vote [169]. Vote-Based [169]. Voting [152, 113].

Weakly [185, 219]. Weakly-Supervised [219]. Web [192, 42, 146, 50, 9, 138]. Web-based [50]. Weighted [169]. why-Questions [119]. Wide [146, 197]. Wide-Coverage [197]. Wikipedia [220]. Wise [137]. without [74]. Word [149, 137, 128, 206, 218, 209, 201, 157, 180, 133, 224, 188, 150, 84, 3, 15, 10, 227, 108]. word-based [15, 10, 108]. Word-Wise [137]. wordnet [6]. Words [187, 116, 170, 203, 93]. World [146]. Writing [214]. WWW [70].

Zero [111]. Zero-anaphora [111].

References

Wong:2002:P

[1] Kam-Fai Wong and Jun’ichi Tsujii. Prologue. ACM Transactions on Asian Language Information Processing, 1(1):1–2, March 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Gao:2002:TUA

[2] Jianfeng Gao, Joshua Goodman, Mingjing Li, and Kai-Fu Lee. Toward a unified approach to statistical language modeling for Chinese. ACM Transactions on Asian Language Information Processing, 1(1):3– 33, March 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lai:2002:MTE

[3] Yu-Sheng Lai and Chung-Hsien Wu. Meaningful term extraction and discriminative term selection in text categorization via unknown-word methodology. ACM Transactions on Asian Language Information Processing, 1(1):34–64, March 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kim:2002:MBG

[4] Byeongchang Kim, Gary Geunbae Lee, and Jong-Hyeok Lee. Morphemebased grapheme to phoneme conversion using phonetic patterns and morphophonemic connectivity information. ACM Transactions on Asian Language Information Processing, 1(1):65–82, March 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lee:2002:UTI

[5] Tan Lee, Wai Lau, Y. W. Wong, and P. C. Ching. Using tone information in Cantonese continuous speech recognition. ACM Transactions on Asian Language Information Processing, 1(1):83–102, March 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2002:BCE

[6] Hsin-Hsi Chen, Chi-Ching Lin, and Wen-Cheng Lin. Building a Chinese– English wordnet for translingual applications. ACM Transactions on Asian Language Information Processing, 1(2):103–122, June 2002. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Meng:2002:GPM

[7] Helen Meng, Po-Chui Luk, Kui Xu, and Fuliang Weng. GLR parsing with multiple grammars for natural language queries. ACM Transactions on Asian Language Information Processing, 1(2):123–144, June 2002. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Murata:2002:CTM

[8] Masaki Murata, Qing Ma, and Hitoshi Isahara. Comparison of three machine-learning methods for Thai part-of-speech tagging. ACM Transactions on Asian Language Information Processing, 1(2):145–158, June 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lu:2002:TWQ

[9] Wen-Hsiang Lu, Lee-Feng Chien, and Hsi-Jian Lee. Translation of Web queries using anchor text mining. ACM Transactions on Asian Language Information Processing, 1(2):159–172, June 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Li:2002:WBA

[10] Wenjie Li and Kam-Fai Wong. A word-based approach for modeling and discovering temporal relations embedded in Chinese sentences. ACM Transactions on Asian Language Information Processing, 1(3):173–206, September 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lee:2002:ACB

[11] Jin-Seok Lee, Byeongchang Kim, and Gary Geunbae Lee. Automatic corpus-based tone and break-index prediction using K-ToBI representation. ACM Transactions on Asian Language Information Processing, 1 (3):207–224, September 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Luk:2002:CCD

[12] Robert W. P. Luk and K. L. Kwok. A comparison of Chinese document indexing strategies and retrieval models. ACM Transactions on Asian Language Information Processing, 1(3):225–268, September 2002. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Suzuki:2002:LCS

[13] Izumi Suzuki, Yoshiki Mikami, Ario Ohsato, and Yoshihide Chubachi. A language and character set determination method based on N-gram statistics. ACM Transactions on Asian Language Information Processing, 1(3):269–278, September 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Jin:2002:CDC

[14] Honglan Jin and Kam-Fai Wong. A Chinese dictionary construction algorithm for information retrieval. ACM Transactions on Asian Language Information Processing, 1(4):281–296, December 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Li:2002:CCB

[15] Yuanxiang Li, Xiaoqing Ding, and Chew Lim Tan. Combining characterbased bigrams with word-based bigrams in contextual postprocessing for Chinese script recognition. ACM Transactions on Asian Language Information Processing, 1(4):297–309, December 2002. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lo:2003:CLS

[16] Wai-Kit Lo, Helen Meng, and P. C. Ching. Cross-language spoken document retrieval using HMM-based retrieval model with multi-scale fusion. ACM Transactions on Asian Language Information Processing, 2 (1):1–26, March 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Shi:2003:OHC

[17] Daming Shi, Robert I. Damper, and Steve R. Gunn. Offline handwritten Chinese character recognition by radical decomposition. ACM Transactions on Asian Language Information Processing, 2(1):27–48, March 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lee:2003:TAS

[18] Yue-Shi Lee. Task adaptation in stochastic language model for Chinese homophone disambiguation. ACM Transactions on Asian Language Information Processing, 2(1):49–62, March 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Shieh:2003:EAT

[19] Jiann-Cherng Shieh. An efficient accessing technique for Taiwanese phonetic transcriptions. ACM Transactions on Asian Language Information Processing, 2(1):63–77, March 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Oard:2003:SLE

[20] Douglas W. Oard. The surprise language exercises. ACM Transactions on Asian Language Information Processing, 2(2):79–84, June 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Allan:2003:MTD

[21] James Allan, Victor Lavrenko, and Margaret E. Connell. A month to topic detection and tracking in Hindi. ACM Transactions on Asian Language Information Processing, 2(2):85–100, June 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Strassel:2003:LRC

[22] Stephanie Strassel, Mike Maxwell, and Christopher Cieri. Linguistic resource creation for research and technology development: a recent experiment. ACM Transactions on Asian Language Information Processing, 2 (2):101–117, June 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Dorr:2003:RPD

[23] Bonnie J. Dorr, Necip Fazil Ayan, Nizar Habash, Nitin Madnani, and Rebecca Hwa. Rapid porting of DUSTer to Hindi. ACM Transactions on Asian Language Information Processing, 2(2):118–123, June 2003. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Huang:2003:ENE

[24] Fei Huang, Stephan Vogel, and Alex Waibel. Extracting named entity translingual equivalence with limited resources. ACM Transactions on Asian Language Information Processing, 2(2):124–129, June 2003. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Larkey:2003:HCT

[25] Leah S. Larkey, Margaret E. Connell, and Nasreen Abduljaleel. Hindi CLIR in thirty days. ACM Transactions on Asian Language Information Processing, 2(2):130–142, June 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lavie:2003:EHE

[26] Alon Lavie, Stephan Vogel, Lori Levin, Erik Peterson, Katharina Probst, Ariadna Font Llitjós, Rachel Reynolds, Jaime Carbonell, and Richard Cohen. Experiments with a Hindi-to-English transfer-based MT system under a miserly data scenario. ACM Transactions on Asian Language Information Processing, 2(2):143–163, June 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Xu:2003:CLR

[27] Jinxi Xu and Ralph Weischedel. Cross-lingual retrieval for Hindi. ACM Transactions on Asian Language Information Processing, 2(2):164–168, June 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

May:2003:SWC

[28] Jonathan May, Ada Brunstein, Prem Natarajan, and Ralph Weischedel. Surprise! What’s in a Cebuano or Hindi Name? ACM Transactions on Asian Language Information Processing, 2(3):169–180, September 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sekine:2003:HEC

[29] Satoshi Sekine and Ralph Grishman. Hindi-English cross-lingual questionanswering system. ACM Transactions on Asian Language Information Processing, 2(3):181–192, September 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Ma:2003:AHO

[30] Huanfeng Ma and David Doermann. Adaptive Hindi OCR using generalized Hausdorff image comparison. ACM Transactions on Asian Language Information Processing, 2(3):193–218, September 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

He:2003:MMI

[31] Daqing He, Douglas W. Oard, Jianqiang Wang, Jun Luo, Dina Demner-Fushman, Kareem Darwish, Philip Resnik, Sanjeev Khudanpur, Michael Nossal, Michael Subotin, and Anton Leuski. Making MIRACLEs: Interactive translingual search for Cebuano and Hindi. ACM Transactions on Asian Language Information Processing, 2(3):219–244, September 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Leuski:2003:CLC

[32] Anton Leuski, Chin-Yew Lin, Liang Zhou, Ulrich Germann, Franz Josef Och, and Eduard Hovy. Cross-lingual C*ST*RD: English access to Hindi information. ACM Transactions on Asian Language Information Processing, 2(3):245–269, September 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Dorr:2003:CLH

[33] Bonnie Dorr, David Zajic, and Richard Schwartz. Cross-language headline generation for Hindi. ACM Transactions on Asian Language Information Processing, 2(3):270–289, September 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Li:2003:RDH

[34] Wei Li and Andrew McCallum. Rapid development of Hindi named entity recognition using conditional random fields and feature induction. ACM Transactions on Asian Language Information Processing, 2(3):290–294, September 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Maynard:2003:RCI

[35] Diana Maynard, Valentin Tablan, Kalina Bontcheva, and Hamish Cunningham. Rapid customization of an information extraction system for a surprise language. ACM Transactions on Asian Language Information Processing, 2(3):295–300, September 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kang:2003:IPP

[36] Mi-Young Kang, Aesun Yoon, and Hyuk-Chul Kwon. Improving partial parsing based on error-pattern analysis for a Korean grammar-checker. ACM Transactions on Asian Language Information Processing, 2(4):301– 323, December 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kim:2003:RRE

[37] Harksoo Kim and Jungyun Seo. Resolution of referring expressions in a Korean multimodal dialogue system. ACM Transactions on Asian Language Information Processing, 2(4):324–337, December 2003. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Mani:2004:ISI

[38] Inderjeet Mani, James Pustejovsky, and Beth Sundheim. Introduction to the special issue on temporal information processing. ACM Transactions on Asian Language Information Processing, 3(1):1–10, March 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Han:2004:FRT

[39] Benjamin Han and Alon Lavie. A framework for resolution of time in natural language. ACM Transactions on Asian Language Information Processing, 3(1):11–32, March 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Schilder:2004:EMT

[40] Frank Schilder. Extracting meaning from temporal nouns and temporal prepositions. ACM Transactions on Asian Language Information Processing, 3(1):33–50, March 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Jang:2004:ATT

[41] Seok Bae Jang, Jennifer Baldwin, and Inderjeet Mani. Automatic TIMEX2 tagging of Korean news. ACM Transactions on Asian Language Information Processing, 3(1):51–65, March 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Hobbs:2004:OTS

[42] Jerry R. Hobbs and Feng Pan. An ontology of time for the Semantic Web. ACM Transactions on Asian Language Information Processing, 3 (1):66–85, March 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Gao:2004:ISI

[43] Jianfeng Gao and Chin-Yew Lin. Introduction to the special issue on statistical language modeling. ACM Transactions on Asian Language Information Processing, 3(2):87–93, June 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kim:2004:LTL

[44] Woosung Kim and Sanjeev Khudanpur. Lexical triggers and latent semantic analysis for cross-lingual language model adaptation. ACM Transactions on Asian Language Information Processing, 3(2):94–112, June 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Linares:2004:HLM

[45] Diego Linares, José-Miguel Benedí, and Joan-Andreu Sánchez. A hybrid language model based on a combination of N-grams and stochastic context-free grammars. ACM Transactions on Asian Language Information Processing, 3(2):113–127, June 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2004:DHG

[46] Berlin Chen, Hsin-Min Wang, and Lin-Shan Lee. A discriminative HMM/ N-gram-based retrieval approach for Mandarin spoken documents. ACM Transactions on Asian Language Information Processing, 3(2):128–145, June 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Nguyen:2004:EBS

[47] Minh Le Nguyen, Susumu Horiguchi, Akira Shimazu, and Bao Tu Ho. Example-based sentence reduction using the hidden Markov model. ACM Transactions on Asian Language Information Processing, 3(2):146–158, June 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Fung:2004:MEC

[48] Pascale Fung, Grace Ngai, Yongsheng Yang, and Benfeng Chen. A maximum-entropy Chinese parser augmented by transformation-based learning. ACM Transactions on Asian Language Information Processing, 3(2):159–168, June 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Li:2004:AMF

[49] Yujia Li, Tan Lee, and Yao Qian. Analysis and modeling of F0 contours for Cantonese text-to-speech. ACM Transactions on Asian Language Information Processing, 3(3):169–180, September 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Huang:2004:UWB

[50] Chien-Chung Huang, Shui-Lung Chuang, and Lee-Feng Chien. Using a Web-based categorization approach to generate thematic metadata from texts. ACM Transactions on Asian Language Information Processing, 3 (3):190–212, September 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Myaeng:2004:ISI

[51] Sung Hyon Myaeng. Introduction to the special issue on computer processing of oriental languages. ACM Transactions on Asian Language Information Processing, 3(4):213, December 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Baoli:2004:ANN

[52] Li Baoli, Lu Qin, and Yu Shiwen. An adaptive k-nearest neighbor text categorization strategy. ACM Transactions on Asian Language Information Processing, 3(4):215–226, December 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kim:2004:UTI

[53] Pyung Kim and Sung Hyon Myaeng. Usefulness of temporal information automatically extracted from news articles for topic tracking. ACM Transactions on Asian Language Information Processing, 3(4):227–242, December 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhang:2004:ESS

[54] Le Zhang, Jingbo Zhu, and Tianshun Yao. An evaluation of statistical spam filtering techniques. ACM Transactions on Asian Language Information Processing, 3(4):243–269, December 2004. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Wu:2005:DSF

[55] Chung-Hsien Wu, Jui-Feng Yeh, and Ming-Jun Chen. Domain-specific FAQ retrieval using independent aspects. ACM Transactions on Asian Language Information Processing, 4(1):1–17, March 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Murata:2005:CEV

[56] Masaki Murata, Masao Utiyama, Kiyotaka Uchimoto, Hitoshi Isahara, and Qing Ma. Correction of errors in a verb modality corpus for machine translation with a machine-learning method. ACM Transactions on Asian Language Information Processing, 4(1):18–37, March 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Hendessi:2005:SSP

[57] F. Hendessi, A. Ghayoori, and T. A. Gulliver. A speech synthesizer for Persian text using a neural network with a smooth ergodic HMM. ACM Transactions on Asian Language Information Processing, 4(1):38– 52, March 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhang:2005:COT

[58] Ying Zhang, Phil Vines, and Justin Zobel. Chinese OOV translation and post-translation query expansion in Chinese–English cross-lingual information retrieval. ACM Transactions on Asian Language Information Processing, 4(2):57–77, June 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Qu:2005:TES

[59] Yan Qu, David A. Hull, Gregory Grefenstette, David A. Evans, Motoko Ishikawa, Setsuko Nara, Toshiya Ueda, Daisuke Noda, Kousaku Arita, Yuki Funakoshi, and Hiroshi Matsuda. Towards effective strategies for monolingual and bilingual information retrieval: Lessons learned from NTCIR-4. ACM Transactions on Asian Language Information Processing, 4(2):78–110, June 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sakai:2005:FPR

[60] Tetsuya Sakai, Toshihiko Manabe, and Makoto Koyama. Flexible pseudorelevance feedback via selective sampling. ACM Transactions on Asian Language Information Processing, 4(2):111–135, June 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kwok:2005:RRP

[61] Kui Lam Kwok, Sora Choi, and Norbert Dinstl. Rich results from poor resources: NTCIR-4 monolingual and cross-lingual retrieval of Korean texts using Chinese and English. ACM Transactions on Asian Language Information Processing, 4(2):135–158, June 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Savoy:2005:CSM

[62] Jacques Savoy. Comparative study of monolingual and multilingual search models for use with Asian languages. ACM Transactions on Asian Language Information Processing, 4(2):159–185, June 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Mase:2005:PTS

[63] Hisao Mase, Tadataka Matsubayashi, Yuichi Ogawa, Makoto Iwayama, and Tadaaki Oshio. Proposal of two-stage patent retrieval method considering the claim structure. ACM Transactions on Asian Language Information Processing, 4(2):186–202, June 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Nakagawa:2005:PSI

[64] Hiroshi Nakagawa, Tatsunori Mori, and Noriko Kando. Preface to the special issues on NTCIR-4. ACM Transactions on Asian Language Information Processing, 4(3):237–242, September 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kato:2005:ODQ

[65] Tsuneaki Kato, Jun’ichi Fukumoto, Fumito Masui, and Noriko Kando. Are open-domain question answering technologies useful for information access dialogues?—an empirical study and a proposal of a novel challenge. ACM Transactions on Asian Language Information Processing, 4(3):243– 262, September 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Isozaki:2005:AHP

[66] Hideki Isozaki. An analysis of a high-performance Japanese question answering system. ACM Transactions on Asian Language Information Processing, 4(3):263–279, September 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Mori:2005:JQA

[67] Tatsunori Mori. Japanese question-answering system using A* search and its improvement. ACM Transactions on Asian Language Information Processing, 4(3):280–304, September 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Mori:2005:MAF

[68] Tatsunori Mori, Masanori Nozawa, and Yoshiaki Asada. Multi-answer-focused multi-document summarization using a question-answering engine. ACM Transactions on Asian Language Information Processing, 4 (3):305–320, September 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Okazaki:2005:ICO

[69] Naoaki Okazaki, Yutaka Matsuo, and Mitsuru Ishizuka. Improving chronological ordering of sentences extracted from multiple newspaper articles. ACM Transactions on Asian Language Information Processing, 4(3):321–339, September 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Yoshioka:2005:CPB

[70] Masaharu Yoshioka and Makoto Haraguchi. On a combination of probabilistic and Boolean IR models for WWW document retrieval. ACM Transactions on Asian Language Information Processing, 4(3):340–356, September 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lingpeng:2005:CIR

[71] Yang Lingpeng, Ji Donghong, Tang Li, and Niu Zhengyu. Chinese information retrieval based on terms and relevant terms. ACM Transactions on Asian Language Information Processing, 4(3):357–374, September 2005. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sakai:2006:ISI

[72] Tetsuya Sakai and Yuji Matsumoto. Introduction to the special issue: Recent advances in information processing and access for Japanese. ACM Transactions on Asian Language Information Processing, 4(4):375–376, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Doi:2006:EBM

[73] Takao Doi, Hirofumi Yamamoto, and Eiichiro Sumita. Example-based machine translation using efficient sentence retrieval based on edit-distance. ACM Transactions on Asian Language Information Processing, 4(4):377– 399, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Tomiura:2006:ESS

[74] Yoichi Tomiura, Shosaku Tanaka, and Toru Hitaka. Estimating satisfactoriness of selectional restriction from corpus without a thesaurus. ACM Transactions on Asian Language Information Processing, 4(4):400–416, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Iida:2006:ARA

[75] Ryu Iida, Kentaro Inui, and Yuji Matsumoto. Anaphora resolution by antecedent identification followed by anaphoricity determination. ACM Transactions on Asian Language Information Processing, 4(4):417–434, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Inui:2006:ACK

[76] Takashi Inui, Kentaro Inui, and Yuji Matsumoto. Acquiring causal knowledge from text using the connective marker tame. ACM Transactions on Asian Language Information Processing, 4(4):435–474, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Ma:2006:TSB

[77] Qiang Ma and Katsumi Tanaka. Topic-structure-based complementary information retrieval and its application. ACM Transactions on Asian Language Information Processing, 4(4):475–503, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Park:2006:ATM

[78] Jong C. Park, Gary Geunbae Lee, and Limsoon Wong. AUTHOR: Text mining and management in biomedicine. ACM Transactions on Asian Language Information Processing, 5(1):1–3, March 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Park:2006:MBB

[79] Kyung-Mi Park, Seon-Ho Kim, Hae-Chang Rim, and Young-Sook Hwang. ME-based biomedical named entity recognition using lexical knowledge. ACM Transactions on Asian Language Information Processing, 5(1):4– 21, March 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Nenadic:2006:MSR

[80] Goran Nenadić and Sophia Ananiadou. Mining semantically related terms from biomedical literature. ACM Transactions on Asian Language Information Processing, 5(1):22–43, March 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kim:2006:ECI

[81] Jung-Jae Kim and Jong C. Park. Extracting contrastive information from negation patterns in biomedical literature. ACM Transactions on Asian Language Information Processing, 5(1):44–60, March 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kim:2006:TPL

[82] Eunju Kim, Yu Song, Cheongjae Lee, Kyoungduk Kim, Gary Geunbae Lee, Byoung-Kee Yi, and Jeongwon Cha. Two-phase learning for biological event extraction and verification. ACM Transactions on Asian Language Information Processing, 5(1):61–73, March 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Mima:2006:TBK

[83] Hideki Mima, Sophia Ananiadou, and Katsumori Matsushima. Terminologybased knowledge mining for new knowledge discovery. ACM Transactions on Asian Language Information Processing, 5(1):74–88, March 2006. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Carpuat:2006:AWS

[84] Marine Carpuat, Pascale Fung, and Grace Ngai. Aligning word senses using bilingual corpora. ACM Transactions on Asian Language Information Processing, 5(2):89–120, June 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lee:2006:ABN

[85] Chun-Jen Lee, Jason S. Chang, and Jyh-Shing R. Jang. Alignment of bilingual named entities in parallel corpora using statistical models and multiple knowledge sources. ACM Transactions on Asian Language Information Processing, 5(2):121–145, June 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Shirado:2006:UJH

[86] Tamotsu Shirado, Satoko Marumoto, Masaki Murata, and Hitoshi Isahara. Using Japanese honorific expressions: a psychological study. ACM Transactions on Asian Language Information Processing, 5(2):146–164, June 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Wu:2006:ERT

[87] Chung-Hsien Wu, Ze-Jing Chuang, and Yu-Chung Lin. Emotion recognition from text using semantic labels and separable mixture models. ACM Transactions on Asian Language Information Processing, 5(2):165–183, June 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Dale:2006:ISS

[88] Robert Dale. Introduction to the Special section: Extended best papers from IJCNLP 2005. ACM Transactions on Asian Language Information Processing, 5(3):183–184, September 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Oh:2006:MTM

[89] Jong-Hoon Oh, Key-Sun Choi, and Hitoshi Isahara. A machine transliteration model based on correspondence between graphemes and phonemes. ACM Transactions on Asian Language Information Processing, 5(3):185– 208, September 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Gao:2006:ESL

[90] Jianfeng Gao, Hisami Suzuki, and Wei Yuan. An empirical study on language model adaptation. ACM Transactions on Asian Language Information Processing, 5(3):209–227, September 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Ye:2006:SRL

[91] Patrick Ye and Timothy Baldwin. Semantic role labeling of prepositional phrases. ACM Transactions on Asian Language Information Processing, 5(3):228–244, September 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chung:2006:APD

[92] Tze Leung Chung, Robert Wing Pong Luk, Kam Fai Wong, Kui Lam Kwok, and Dik Lun Lee. Adapting pivoted document-length normalization for query size: Experiments in Chinese and English. ACM Transactions on Asian Language Information Processing, 5(3):245–263, September 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Matsumura:2006:ERB

[93] Atsushi Matsumura, Atsuhiro Takasu, and Jun Adachi. Effect of relationships between words on Japanese information retrieval. ACM Transactions on Asian Language Information Processing, 5(3):264–289, September 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Song:2006:ISI

[94] Dawei Song and Jian-Yun Nie. Introduction to special issue on reasoning in natural language information processing. ACM Transactions on Asian Language Information Processing, 5(4):291–295, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Nie:2006:ILM

[95] Jian-Yun Nie, Guihong Cao, and Jing Bai. Inferential language models for information retrieval. ACM Transactions on Asian Language Information Processing, 5(4):296–322, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Gao:2006:SQT

[96] Jianfeng Gao, Jian-Yun Nie, and Ming Zhou. Statistical query translation models for cross-language information retrieval. ACM Transactions on Asian Language Information Processing, 5(4):323–359, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Liu:2006:SFQ

[97] Yi Liu, Rong Jin, and Joyce Y. Chai. A statistical framework for query translation disambiguation. ACM Transactions on Asian Language Information Processing, 5(4):360–387, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Li:2006:TTT

[98] Baoli Li, Wenjie Li, and Qin Lu. Topic tracking with time granularity reasoning. ACM Transactions on Asian Language Information Processing, 5(4):388–412, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Phan:2006:IDS

[99] Xuan-Hieu Phan, Le-Minh Nguyen, Yasushi Inoguchi, Tu-Bao Ho, and Susumu Horiguchi. Improving discriminative sequential learning by discovering important association of statistics. ACM Transactions on Asian Language Information Processing, 5(4):413–438, December 2006. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2007:UDM

[100] Yong Chen and Kwok-Ping Chan. Using data mining techniques and rough set theory for language modeling. ACM Transactions on Asian Language Information Processing, 6(1):??, April 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Hsu:2007:MSB

[101] Chung-Chian Hsu, Chien-Hsing Chen, Tien-Teng Shih, and Chun-Kai Chen. Measuring similarity between transliterations against noise data. ACM Transactions on Asian Language Information Processing, 6(1):??, April 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sakai:2007:RFQ

[102] Tetsuya Sakai. On the reliability of factoid question answering evaluation. ACM Transactions on Asian Language Information Processing, 6(1):??, April 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Wiseman:2007:CBC

[103] Yair Wiseman and Irit Gefner. Conjugation-based compression for Hebrew texts. ACM Transactions on Asian Language Information Processing, 6 (1):??, April 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Wu:2007:TBS

[104] Chung-Hsien Wu, Hung-Yu Su, Yu-Hsien Chiu, and Chia-Hung Lin. Transfer-based statistical translation of Taiwanese sign language using PCFG. ACM Transactions on Asian Language Information Processing, 6(1):??, April 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kuo:2007:PSM

[105] Jin-Shea Kuo, Haizhou Li, and Ying-Kuei Yang. A phonetic similarity model for automatic extraction of transliteration pairs. ACM Transactions on Asian Language Information Processing, 6(2):6:1–6:??, September 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Xiao:2007:SNM

[106] Jinghui Xiao, Xiaolong Wang, and Bingquan Liu. The study of a nonstationary maximum entropy Markov model and its application on the pos-tagging task. ACM Transactions on Asian Language Information Processing, 6(2):7:1–7:??, September 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhuang:2007:IHD

[107] Yl Zhuang, Yueting Zhuang, Qing Li, and Lei Chen. Interactive highdimensional index for large Chinese calligraphic character databases. ACM Transactions on Asian Language Information Processing, 6(2):8:1–8:??, September 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Saraswathi:2007:CPE

[108] S. Saraswathi and T. V. Geetha. Comparison of performance of enhanced morpheme-based language model with different word-based language models for improving the performance of Tamil speech recognition system. ACM Transactions on Asian Language Information Processing, 6(3):9:1– 9:??, November 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Hussain:2007:DLS

[109] Sarmad Hussain, Sana Gul, and Afifah Waseem. Developing lexicographic sorting: An example for Urdu. ACM Transactions on Asian Language Information Processing, 6(3):10:1–10:??, November 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Fukumoto:2007:TTB

[110] Fumiyo Fukumoto and Yoshimi Suzuki. Topic tracking based on bilingual comparable corpora and semisupervised clustering. ACM Transactions on Asian Language Information Processing, 6(3):11:1–11:??, November 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Iida:2007:ZAR

[111] Ryu Iida, Kentaro Inui, and Yuji Matsumoto. Zero-anaphora resolution by learning rich syntactic pattern features. ACM Transactions on Asian Language Information Processing, 6(4):1:1–1:22, December 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Adriani:2007:SIC

[112] Mirna Adriani, Jelita Asian, Bobby Nazief, S. M. M. Tahaghoghi, and Hugh E. Williams. Stemming Indonesian: a confix-stripping approach. ACM Transactions on Asian Language Information Processing, 6(4):2:1– 2:33, December 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Thao:2007:NER

[113] Pham Thi Xuan Thao, Tran Quoc Tri, Dinh Dien, and Nigel Collier. Named entity recognition in Vietnamese using classifier voting. ACM Transactions on Asian Language Information Processing, 6(4):3:1–3:18, December 2007. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2008:SBM

[114] Yufeng Chen and Chengqing Zong. A structure-based model for Chinese organization name translation. ACM Transactions on Asian Language Information Processing, 7(1):1:1–1:??, February 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Jeong:2008:ISR

[115] Minwoo Jeong and Gary Geunbae Lee. Improving speech recognition and understanding using error-corrective reranking. ACM Transactions on Asian Language Information Processing, 7(1):2:1–2:??, February 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kuo:2008:MSG

[116] June-Jei Kuo and Hsin-Hsi Chen. Multidocument summary generation: Using informative and event words. ACM Transactions on Asian Language Information Processing, 7(1):3:1–3:??, February 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kando:2008:INS

[117] Noriko Kando, Teruko Mitamura, and Tetsuya Sakai. Introduction to the NTCIR-6 Special Issue. ACM Transactions on Asian Language Information Processing, 7(2):4:1–4:??, June 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhou:2008:HTE

[118] Dong Zhou, Mark Truran, Tim Brailsford, and Helen Ashman. A hybrid technique for English–Chinese cross language information retrieval. ACM Transactions on Asian Language Information Processing, 7(2):5:1–5:??, June 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Higashinaka:2008:AAC

[119] Ryuichiro Higashinaka and Hideki Isozaki. Automatically acquiring causal expression patterns from relation-annotated corpora to improve question answering for why-questions. ACM Transactions on Asian Language Information Processing, 7(2):6:1–6:??, June 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Li:2008:ASV

[120] Yaoyong Li and Kalina Bontcheva. Adapting support vector machines for F -term-based classification of patents. ACM Transactions on Asian Language Information Processing, 7(2):7:1–7:??, June 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Fukumoto:2008:ICL

[121] Fumiyo Fukumoto and Yoshimi Suzuki. Integrating cross-language hierarchies and its application to retrieving relevant documents. ACM Transactions on Asian Language Information Processing, 7(3):8:1–8:??, August 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sharma:2008:AMI

[122] Utpal Sharma, Jugal K. Kalita, and Rajib K. Das. Acquisition of morphology of an Indic language from text corpus. ACM Transactions on Asian Language Information Processing, 7(3):9:1–9:??, August 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2008:TTR

[123] Jiang-Chun Chen and Jyh-Shing Roger Jang. TRUES: Tone Recognition Using Extended Segments. ACM Transactions on Asian Language Information Processing, 7(3):10:1–10:??, August 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lin:2008:VCD

[124] Jeng-Wei Lin, Jan-Ming Ho, Li-Ming Tseng, and Feipei Lai. Variant Chinese domain name resolution. ACM Transactions on Asian Language Information Processing, 7(4):11:1–11:??, November 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lee:2008:BCQ

[125] Cheng-Wei Lee, Min-Yuh Day, Cheng-Lung Sung, Yi-Hsun Lee, Tian-Jian Jiang, Chia-Wei Wu, Cheng-Wei Shih, Yu-Ren Chen, and Wen-Lian Hsu. Boosting Chinese question answering with two lightweight methods: ABSPs and SCO-QAT. ACM Transactions on Asian Language Information Processing, 7(4):12:1–12:??, November 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Che:2008:UHC

[126] Wanxiang Che, Min Zhang, AiTi Aw, ChewLim Tan, Ting Liu, and Sheng Li. Using a hybrid convolution tree kernel for semantic role labeling. ACM Transactions on Asian Language Information Processing, 7(4):13:1– 13:??, November 2008. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Wu:2009:ISI

[127] Chung-Hsien Wu and Haizhou Li. Introduction to the special issue on recent advances in Asian language spoken document retrieval. ACM Transactions on Asian Language Information Processing, 8(1):1:1–1:??, March 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2009:WTM

[128] Berlin Chen. Word topic models for spoken document retrieval and transcription. ACM Transactions on Asian Language Information Processing, 8(1):2:1–2:??, March 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lin:2009:CSP

[129] Shih-Hsiang Lin, Berlin Chen, and Hsin-Min Wang. A comparative study of probabilistic ranking models for Chinese spoken document summarization. ACM Transactions on Asian Language Information Processing, 8(1): 3:1–3:??, March 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2009:TSH

[130] Boxing Chen, Min Zhang, and Ai Ti Aw. Two-stage hypotheses generation for spoken language translation. ACM Transactions on Asian Language Information Processing, 8(1):4:1–4:??, March 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chiang:2009:ISI

[131] David Chiang and Philipp Koehn. Introduction to the special issue on machine translation of Asian language. ACM Transactions on Asian Language Information Processing, 8(2):5:1–5:??, May 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

He:2009:IMH

[132] Xiaodong He, Mei Yang, Jianfeng Gao, Patrick Nguyen, and Robert Moore. Improved monolingual hypothesis alignment for machine translation system combination. ACM Transactions on Asian Language Information Processing, 8(2):6:1–6:??, May 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Ma:2009:BMW

[133] Yanjun Ma and Andy Way. Bilingually motivated word segmentation for statistical machine translation. ACM Transactions on Asian Language Information Processing, 8(2):7:1–7:??, May 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Venkatapathy:2009:DMT

[134] Sriram Venkatapathy and Srinivas Bangalore. Discriminative machine translation using global lexical selection. ACM Transactions on Asian Language Information Processing, 8(2):8:1–8:??, May 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Tsunakawa:2009:CJL

[135] Takashi Tsunakawa, Naoaki Okazaki, Xiao Liu, and Jun’ichi Tsujii. A Chinese–Japanese lexical machine translation through a pivot language. ACM Transactions on Asian Language Information Processing, 8(2):9:1– 9:??, May 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2009:USD

[136] Wenliang Chen, Daisuke Kawahara, Kiyotaka Uchimoto, Yujie Zhang, and Hitoshi Isahara. Using short dependency relations from auto-parsed data for Chinese dependency parsing. ACM Transactions on Asian Language Information Processing, 8(3):10:1–10:??, August 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chanda:2009:WWT

[137] Sukalpa Chanda, Umapada Pal, and Oriol Ramos Terrades. Word-wise Thai and Roman script identification. ACM Transactions on Asian Language Information Processing, 8(3):11:1–11:??, August 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Nguyen:2009:WSC

[138] Cam-Tu Nguyen, Xuan-Hieu Phan, Susumu Horiguchi, Thu-Trang Nguyen, and Quang-Thuy Ha. Web search clustering and labeling with hidden topics. ACM Transactions on Asian Language Information Processing, 8(3):12:1–12:??, August 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Shaalan:2009:ISI

[139] K. Shaalan and A. Farghaly. Introduction to the special issue on Arabic natural language processing. ACM Transactions on Asian Language Information Processing, 8(4):13:1–13:??, December 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Farghaly:2009:ANL

[140] Ali Farghaly and Khaled Shaalan. Arabic natural language processing: Challenges and solutions. ACM Transactions on Asian Language Information Processing, 8(4):14:1–14:??, December 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Espana-Bonet:2009:DPB

[141] Cristina España-Bonet, Jesús Giménez, and Lluís Màrquez. Discriminative phrase-based models for Arabic machine translation. ACM Transactions on Asian Language Information Processing, 8(4):15:1–15:??, December 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Benajiba:2009:MBS

[142] Yassine Benajiba and Imed Zitouni. Morphology-based segmentation combination for Arabic mention detection. ACM Transactions on Asian Language Information Processing, 8(4):16:1–16:??, December 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zitouni:2009:CLI

[143] Imed Zitouni and Radu Florian. Cross-language information propagation for Arabic mention detection. ACM Transactions on Asian Language Information Processing, 8(4):17:1–17:??, December 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lamel:2009:AST

[144] Lori Lamel, Abdelkhalek Messaoudi, and Jean-Luc Gauvain. Automatic speech-to-text transcription in Arabic. ACM Transactions on Asian Language Information Processing, 8(4):18:1–18:??, December 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Moisl:2009:SLL

[145] Hermann Moisl. Sura length and lexical probability estimation in cluster analysis of the Qur’an. ACM Transactions on Asian Language Information Processing, 8(4):19:1–19:??, December 2009. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Hsu:2010:MST

[146] Chung-Chian Hsu and Chien-Hsing Chen. Mining synonymous transliterations from the World Wide Web. ACM Transactions on Asian Language Information Processing, 9(1):1:1–1:??, March 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Liu:2010:ISS

[147] Feifan Liu and Yang Liu. Identification of soundbite and its speaker name using transcripts of broadcast news speech. ACM Transactions on Asian Language Information Processing, 9(1):2:1–2:??, March 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Tepper:2010:IMU

[148] Michael Tepper and Fei Xia. Inducing morphemes using light knowledge. ACM Transactions on Asian Language Information Processing, 9(1):3:1– 3:??, March 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Baldwin:2010:RMB

[149] Timothy Baldwin, Sunam Kim, Francis Bond, Sanae Fujita, David Martinez, and Takaaki Tanaka. A reexamination of MRD-based word sense disambiguation. ACM Transactions on Asian Language Information Processing, 9(1):4:1–4:??, March 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhao:2010:UCB

[150] Hai Zhao, Chang-Ning Huang, Mu Li, and Bao-Liang Lu. A unified character-based tagging framework for Chinese word segmentation. ACM Transactions on Asian Language Information Processing, 9(2):5:1–5:??, June 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Guo:2010:LIS

[151] Yuqing Guo, Haifeng Wang, and Josef van Genabith. A linguistically inspired statistical model for Chinese punctuation generation. ACM Transactions on Asian Language Information Processing, 9(2):6:1–6:??, June 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Naptali:2010:TDL

[152] Welly Naptali, Masatoshi Tsuchiya, and Seiichi Nakagawa. Topicdependent language model with voting on noun history. ACM Transactions on Asian Language Information Processing, 9(2):7:1–7:??, June 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Ng:2010:SJ

[153] Hwee Tou Ng. The state of the journal. ACM Transactions on Asian Language Information Processing, 9(3):8:1–8:??, September 2010. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Harman:2010:ISI

[154] Donna Harman, Noriko Kando, Prasenjit Majumder, Mandar Mitra, and Carol Peters. Introduction to the Special Issue on Indian Language Information Retrieval Part I. ACM Transactions on Asian Language Information Processing, 9(3):9:1–9:??, September 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Majumder:2010:FEE

[155] Prasenjit Majumder, Mandar Mitra, Dipasree Pal, Ayan Bandyopadhyay, Samaresh Maiti, Sukomal Pal, Deboshree Modak, and Sucharita Sanyal. The FIRE 2008 evaluation exercise. ACM Transactions on Asian Language Information Processing, 9(3):10:1–10:??, September 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Dolamic:2010:CSI

[156] Ljiljana Dolamic and Jacques Savoy. Comparative study of indexing and search strategies for the Hindi, Marathi, and Bengali languages. ACM Transactions on Asian Language Information Processing, 9(3):11:1–11:??, September 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Leveling:2010:SWI

[157] Johannes Leveling and Gareth J. F. Jones. Sub-word indexing and blind relevance feedback for English, Bengali, Hindi, and Marathi IR. ACM Transactions on Asian Language Information Processing, 9(3):12:1–12:??, September 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kumaran:2010:CMT

[158] A. Kumaran, Mitesh M. Khapra, and Pushpak Bhattacharyya. Compositional machine transliteration. ACM Transactions on Asian Language Information Processing, 9(4):13:1–13:??, December 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chinnakotla:2010:TRS

[159] Manoj K. Chinnakotla, Om P. Damani, and Avijit Satoskar. Transliteration for resource-scarce languages. ACM Transactions on Asian Language Information Processing, 9(4):14:1–14:??, December 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Mukund:2010:IES

[160] Smruthi Mukund, Rohini Srihari, and Erik Peterson. An informationextraction system for Urdu—A resource-poor language. ACM Transactions on Asian Language Information Processing, 9(4):15:1–15:??, December 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Katz:2011:ISI

[161] Graham Katz and Mona Diab. Introduction to the special issue on Arabic computational linguistics. ACM Transactions on Asian Language Information Processing, 10(1):1:1–1:??, March 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Condon:2011:MTE

[162] S. Condon, D. Parvaz, J. Aberdeen, C. Doran, A. Freeman, and M. Awad. Machine translation errors: English and Iraqi Arabic. ACM Transactions on Asian Language Information Processing, 10(1):2:1–2:??, March 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Rytting:2011:SCD

[163] C. Anton Rytting, David M. Zajic, Paul Rodrigues, Sarah C. Wayland, Christian Hettick, Tim Buckwalter, and Charles C. Blake. Spelling correction for dialectal Arabic dictionary lookup. ACM Transactions on Asian Language Information Processing, 10(1):3:1–3:??, March 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kulick:2011:ESC

[164] Seth Kulick. Exploiting separation of closed-class categories for Arabic tokenization and part-of-speech tagging. ACM Transactions on Asian Language Information Processing, 10(1):4:1–4:??, March 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Abdul-Mageed:2011:ADA

[165] Muhammad Abdul-Mageed. Automatic detection of Arabic non-anaphoric pronouns for improving anaphora resolution. ACM Transactions on Asian Language Information Processing, 10(1):5:1–5:??, March 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Wu:2011:IPD

[166] Chung-Hsien Wu, Wei-Bin Liang, and Jui-Feng Yeh. Interruption point detection of spontaneous speech using inter-syllable boundary-based prosodic features. ACM Transactions on Asian Language Information Processing, 10(1):6:1–6:??, March 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Wu:2011:ADS

[167] Chung-Hsien Wu, Hung-Yu Su, and Han-Ping Shen. Articulationdisordered speech recognition using speaker-adaptive acoustic models and personalized articulation patterns. ACM Transactions on Asian Language Information Processing, 10(2):7:1–7:??, June 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Paik:2011:FCB

[168] Jiaul H. Paik and Swapan K. Parui. A fast corpus-based stemmer. ACM Transactions on Asian Language Information Processing, 10(2):8:1–8:??, June 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Ekbal:2011:WVB

[169] Asif Ekbal and Sriparna Saha. Weighted vote-based classifier ensemble for named entity recognition: a genetic algorithm-based approach. ACM Transactions on Asian Language Information Processing, 10(2):9:1–9:??, June 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Liu:2011:VPS

[170] C.-L. Liu, M.-H. Lai, K.-W. Tien, Y.-H. Chuang, S.-H. Wu, and C.-Y. Lee. Visually and phonologically similar characters in incorrect Chinese words: Analyses, identification, and applications. ACM Transactions on Asian Language Information Processing, 10(2):10:1–10:??, June 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chen:2011:ISI

[171] Keh-Jiann Chen, Qun Liu, Nianwen Xue, and Le Sun. Introduction to the special issue on Chinese language processing. ACM Transactions on Asian Language Information Processing, 10(3):11:1–11:??, September 2011. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhu:2011:ATC

[172] Muhua Zhu, Jingbo Zhu, and Tong Xiao. Automatic treebank conversion via informed decoding — A case study on Chinese treebanks. ACM Transactions on Asian Language Information Processing, 10(3): 12:1–12:??, September 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Li:2011:USR

[173] Junhui Li and Guodong Zhou. Unified semantic role labeling for verbal and nominal predicates in the Chinese language. ACM Transactions on Asian Language Information Processing, 10(3):13:1–13:??, September 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhang:2011:DPS

[174] Peng Zhang, Wenjie Li, Yuexian Hou, and Dawei Song. Developing position structure-based framework for Chinese entity relation extraction.

ACM Transactions on Asian Language Information Processing, 10(3): 14:1–14:??, September 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Qian:2011:ECD

[175] Longhua Qian, Guodong Zhou, and Qiaoming Zhu. Employing constituent dependency information for tree kernel-based semantic relation extraction between named entities. ACM Transactions on Asian Language Information Processing, 10(3):15:1–15:??, September 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Huang:2011:UST

[176] Chung-Chi Huang, Ho-Ching Yen, Ping-Che Yang, Shih-Ting Huang, and Jason S. Chang. Using sublexical translations to handle the OOV problem in machine translation. ACM Transactions on Asian Language Information Processing, 10(3):16:1–16:??, September 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Du:2011:ICE

[177] Jinhua Du and Andy Way. Improved Chinese–English SMT with Chinese “DE’’ construction classification and reordering. ACM Transactions on Asian Language Information Processing, 10(4):17:1–17:??, December 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Xiao:2011:LMS

[178] Tong Xiao, Jingbo Zhu, and Muhua Zhu. Language modeling for Syntax-Based machine translation using tree substitution grammars: a case study on Chinese-English translation. ACM Transactions on Asian Language Information Processing, 10(4):18:1–18:??, December 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Li:2011:MEC

[179] Lishuang Li, Peng Wang, Degen Huang, and Lian Zhao. Mining English– Chinese named entity pairs from comparable corpora. ACM Transactions on Asian Language Information Processing, 10(4):19:1–19:??, December 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Liu:2011:UBR

[180] Zhiyuan Liu, Yabin Zheng, Lixing Xie, Maosong Sun, Liyun Ru, and Yang Zhang. User behaviors in related word retrieval and new word detection: a collaborative perspective. ACM Transactions on Asian Language Information Processing, 10(4):20:1–20:??, December 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Wang:2011:DLA

[181] Baoxun Wang, Bingquan Liu, Xiaolong Wang, Chengjie Sun, and Deyuan Zhang. Deep learning approaches to semantic relevance modeling for Chinese question–answer pairs. ACM Transactions on Asian Language Information Processing, 10(4):21:1–21:??, December 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Pal:2012:HRI

[182] Umapada Pal, Ramachandran Jayadevan, and Nabin Sharma. Handwriting recognition in Indian regional scripts: a survey of offline techniques. ACM Transactions on Asian Language Information Processing, 11(1):1:1– 1:??, March 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zaghouani:2012:RRB

[183] Wajdi Zaghouani. RENAR: a rule-based Arabic named entity recognition system. ACM Transactions on Asian Language Information Processing, 11(1):2:1–2:??, March 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chang:2012:EDC

[184] Ru-Yng Chang, Chung-Hsien Wu, and Philips Kokoh Prasetyo. Error diagnosis of Chinese sentences using inductive learning algorithm and decomposition-based testing mechanism. ACM Transactions on Asian Language Information Processing, 11(1):3:1–3:??, March 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

He:2012:ISP

[185] Yulan He. Incorporating sentiment prior knowledge for weakly supervised sentiment analysis. ACM Transactions on Asian Language Information Processing, 11(2):4:1–4:??, June 2012. ISSN 1530-0226 (print), 1558-3430 (electronic).

Wang:2012:TUF

[186] Hongling Wang and Guodong Zhou. Toward a unified framework for standard and update multi-document summarization. ACM Transactions on Asian Language Information Processing, 11(2):5:1–5:??, June 2012. ISSN 1530-0226 (print), 1558-3430 (electronic).

Andrade:2012:SEC

[187] Daniel Andrade, Takuya Matsuzaki, and Jun’ichi Tsujii. Statistical extraction and comparison of pivot words for bilingual lexicon extension.

ACM Transactions on Asian Language Information Processing, 11(2):6:1– 6:??, June 2012. ISSN 1530-0226 (print), 1558-3430 (electronic).

Wang:2012:IGD

[188] Kun Wang, Chengqing Zong, and Keh-Yih Su. Integrating generative and discriminative character-based models for Chinese word segmentation. ACM Transactions on Asian Language Information Processing, 11(2):7:1– 7:??, June 2012. ISSN 1530-0226 (print), 1558-3430 (electronic).

Isozaki:2012:HBP

[189] Hideki Isozaki, Katsuhito Sudoh, Hajime Tsukada, and Kevin Duh. HPSG-based preprocessing for English-to-Japanese translation. ACM Transactions on Asian Language Information Processing, 11(3):8:1–8:??, September 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhang:2012:ABH

[190] Lidan Zhang and Kwop-Ping Chan. Adaptive Bayesian HMM for fully unsupervised Chinese part-of-speech induction. ACM Transactions on Asian Language Information Processing, 11(3):9:1–9:??, September 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Lee:2012:SMB

[191] Jinsik Lee, Sungjin Lee, Jonghoon Lee, Byeongchang Kim, and Gary Geunbae Lee. Stacking model-based Korean prosodic phrasing using speaker variability reduction and linguistic feature engineering. ACM Transactions on Asian Language Information Processing, 11(3):10:1–10:??, September 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Duc:2012:CLL

[192] Nguyen Tuan Duc, Danushka Bollegala, and Mitsuru Ishizuka. Crosslanguage latent relational search between Japanese and English languages using a Web corpus. ACM Transactions on Asian Language Information Processing, 11(3):11:1–11:??, September 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Mitamura:2012:ISI

[193] Teruko Mitamura, Noriko Kando, and Koichi Takeda. Introduction to the special issue on RITE. ACM Transactions on Asian Language Information Processing, 11(4):12:1–12:??, December 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Miyao:2012:ETE

[194] Yusuke Miyao, Hideki Shima, Hiroshi Kanayama, and Teruko Mitamura. Evaluating textual entailment recognition for university entrance examinations. ACM Transactions on Asian Language Information Processing, 11(4):13:1–13:??, December 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Pham:2012:LRT

[195] Minh Quang Nhat Pham, Minh Le Nguyen, and Akira Shimazu. Learning to recognize textual entailment in Japanese texts with the utilization of machine translation. ACM Transactions on Asian Language Information Processing, 11(4):14:1–14:??, December 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Qiu:2012:RIT

[196] Xipeng Qiu, Ling Cao, Zhao Liu, and Xuanjing Huang. Recognizing inference in texts with Markov logic networks. ACM Transactions on Asian Language Information Processing, 11(4):15:1–15:??, December 2012. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Shibata:2012:PAS

[197] Tomohide Shibata and Sadao Kurohashi. Predicate-argument structurebased textual entailment recognition system exploiting wide-coverage lexical knowledge. ACM Transactions on Asian Language Information Processing, 11(4):16:1–16:??, December 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Shih:2012:VCT

[198] Chengwei Shih, Chengwei Lee, Richard Tzonghan Tsai, and Wenlian Hsu. Validating contradiction in texts using online co-mention pattern checking. ACM Transactions on Asian Language Information Processing, 11 (4):17:1–17:??, December 2012. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Watanabe:2012:LDL

[199] Yotaro Watanabe, Junta Mizuno, Eric Nichols, Katsuma Narisawa, Keita Nabeshima, Naoaki Okazaki, and Kentaro Inui. Leveraging diverse lexical resources for textual entailment recognition. ACM Transactions on Asian Language Information Processing, 11(4):18:1–18:??, December 2012. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Hao:2013:TPP

[200] Tianyong Hao and Chunshen Zhu. Toward a professional platform for Chinese character conversion. ACM Transactions on Asian Language Information Processing, 12(1):1:1–1:??, March 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Jiang:2013:LRC

[201] Mike Tian-Jian Jiang, Tsung-Hsien Lee, and Wen-Lian Hsu. The left and right context of a word: Overlapping Chinese syllable word segmentation with minimal context. ACM Transactions on Asian Language Information Processing, 12(1):2:1–2:??, March 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Bach:2013:TPF

[202] Ngo Xuan Bach, Nguyen Le Minh, Tran Thi Oanh, and Akira Shimazu. A two-phase framework for learning logical structures of paragraphs in legal articles. ACM Transactions on Asian Language Information Processing, 12(1):3:1–3:??, March 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sundaram:2013:AFB

[203] Suresh Sundaram and A. G. Ramakrishnan. Attention-feedback based robust segmentation of online handwritten isolated Tamil words. ACM Transactions on Asian Language Information Processing, 12(1):4:1–4:??, March 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sun:2013:LAC

[204] Xu Sun, Naoaki Okazaki, Jun’ichi Tsujii, and Houfeng Wang. Learning abbreviations from Chinese and English terms by modeling non-local information. ACM Transactions on Asian Language Information Processing, 12(2):5:1–5:??, June 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Hinkle:2013:DES

[205] Lauren Hinkle, Albert Brouillette, Sujay Jayakar, Leigh Gathings, Miguel Lezcano, and Jugal Kalita. Design and evaluation of soft keyboards for Brahmic scripts. ACM Transactions on Asian Language Information Processing, 12(2):6:1–6:??, June 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Fujita:2013:WSD

[206] Sanae Fujita and Akinori Fujino. Word sense disambiguation by combining labeled data expansion and semi-supervised learning method. ACM Transactions on Asian Language Information Processing, 12(2):7:1–7:??, June 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sproat:2013:EGN

[207] Richard Sproat. Editorial greetings from the new Editor-in–Chief. ACM Transactions on Asian Language Information Processing, 12(3):8:1–8:??, August 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Fukunishi:2013:BAA

[208] Takaaki Fukunishi, Andrew Finch, Seiichi Yamamoto, and Eiichiro Sumita. A Bayesian alignment approach to transliteration mining. ACM Transactions on Asian Language Information Processing, 12(3):9:1–9:??, August 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Iwakura:2013:NER

[209] Tomoya Iwakura, Hiroya Takamura, and Manabu Okumura. A named entity recognition method based on decomposition and concatenation of word chunks. ACM Transactions on Asian Language Information Processing, 12(3):10:1–10:??, August 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Izumi:2013:NCF

[210] Tomoko Izumi, Kenji Imamura, Taichi Asami, Kuniko Saito, Genichiro Kikui, and Satoshi Sato. Normalizing complex functional expressions in Japanese predicates: Linguistically-directed rule-based paraphrasing and its application. ACM Transactions on Asian Language Information Processing, 12(3):11:1–11:??, August 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sudoh:2013:SBP

[211] Katsuhito Sudoh, Xianchao Wu, Kevin Duh, Hajime Tsukada, and Masaaki Nagata. Syntax-based post-ordering for efficient Japanese-to-English translation. ACM Transactions on Asian Language Information Processing, 12(3):12:1–12:??, August 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sproat:2013:TP

[212] Richard Sproat. TALIP perspectives. ACM Transactions on Asian Language Information Processing, 12(4):13:1–13:??, October 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Paul:2013:HCB

[213] Michael Paul, Andrew Finch, and Eiichrio Sumita. How to choose the best pivot language for automatic translation of low-resource languages. ACM Transactions on Asian Language Information Processing, 12(4): 14:1–14:??, October 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Huang:2013:CAT

[214] Chung-Chi Huang, Mei-Hua Chen, Ping-Che Yang, and Jason S. Chang. A computer-assisted translation and writing system. ACM Transactions on Asian Language Information Processing, 12(4):15:1–15:??, October 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Chu:2013:CJM

[215] Chenhui Chu, Toshiaki Nakazawa, Daisuke Kawahara, and Sadao Kurohashi. Chinese–Japanese machine translation exploiting Chinese characters. ACM Transactions on Asian Language Information Processing, 12 (4):16:1–16:??, October 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Goto:2013:POP

[216] Isao Goto, Masao Utiyama, and Eiichiro Sumita. Post-ordering by parsing with ITG for Japanese–English statistical machine translation. ACM Transactions on Asian Language Information Processing, 12(4): 17:1–17:??, October 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Strotgen:2014:TML

[217] Jannik Strötgen, Ayser Armiti, Tran Van Canh, Julian Zell, and Michael Gertz. Time for more languages: Temporal tagging of Arabic, Italian, Spanish, and Vietnamese. ACM Transactions on Asian Language Information Processing, 13(1):1:1–1:??, February 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Goto:2014:DMB

[218] Isao Goto, Masao Utiyama, Eiichiro Sumita, Akihiro Tamura, and Sadao Kurohashi. Distortion model based on word sequence labeling for statistical machine translation. ACM Transactions on Asian Language Information Processing, 13(1):2:1–2:??, February 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kim:2014:CLA

[219] Seokhwan Kim, Minwoo Jeong, Jonghoon Lee, and Gary Geunbae Lee. Cross-lingual annotation projection for weakly-supervised relation extraction. ACM Transactions on Asian Language Information Processing, 13 (1):3:1–3:??, February 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Yahya:2014:ATC

[220] Adnan Yahya and Ali Salhi. Arabic text categorization based on Arabic Wikipedia. ACM Transactions on Asian Language Information Processing, 13(1):4:1–4:??, February 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Church:2014:TPG

[221] Kenneth Church. TALIP perspectives, guest editorial commentary: What counts (and what ought to count)? ACM Transactions on Asian Language Information Processing, 13(1):5:1–5:??, February 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sulaiman:2014:EJS

[222] Suliana Sulaiman, Khairuddin Omar, Nazlia Omar, Mohd Zamri Murah, and Hamdan Abdul Rahman. The effectiveness of a Jawi stemmer for retrieving relevant Malay documents in Jawi characters. ACM Transactions on Asian Language Information Processing, 13(2):6:1–6:??, June 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Esmaili:2014:TKI

[223] Kyumars Sheykh Esmaili, Shahin Salavati, and Anwitaman Datta. Towards Kurdish information retrieval. ACM Transactions on Asian Language Information Processing, 13(2):7:1–7:??, June 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sharma:2014:WPS

[224] Manoj Kumar Sharma and Debasis Samanta. Word prediction system for text entry in Hindi. ACM Transactions on Asian Language Information Processing, 13(2):8:1–8:??, June 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Keskes:2014:SAT

[225] Iskandar Keskes, Farah Benamara Zitoune, and Lamia Hadrich Belguith. Splitting Arabic texts into elementary discourse units. ACM Transactions on Asian Language Information Processing, 13(2):9:1–9:??, June 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Rubin:2014:TPG

[226] Victoria L. Rubin. TALIP perspectives, guest editorial commentary: Pragmatic and cultural considerations for deception detection in Asian languages. ACM Transactions on Asian Language Information Processing, 13(2):10:1–10:??, June 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Na:2014:LAN

[227] Hwidong Na and Jong-Hyeok Lee. Linguistic analysis of non-ITG word reordering between language pairs with different word order typologies. ACM Transactions on Asian Language Information Processing, 13(3): 11:1–11:??, September 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

A:2014:AMO

[228] Bharath A. and Sriganesh Madhvanath. Allograph modeling for online handwritten characters in Devanagari using constrained stroke clustering. ACM Transactions on Asian Language Information Processing, 13 (3):12:1–12:??, September 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Paik:2014:IBF

[229] Jiaul H. Paik, Dipasree Pal, and Swapan K. Parui. Incremental blind feedback: an effective approach to automatic query expansion. ACM Transactions on Asian Language Information Processing, 13(3):13:1– 13:??, September 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Saharia:2014:SRP

[230] Navanath Saharia, Utpal Sharma, and Jugal Kalita. Stemming resourcepoor Indian languages. ACM Transactions on Asian Language Information Processing, 13(3):14:1–14:??, September 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sproat:2014:SJ

[231] Richard Sproat. The state of the journal. ACM Transactions on Asian Language Information Processing, 13(3):15:1–15:??, September 2014. CO-DEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Bang:2014:PVP

[232] Jeesoo Bang, Jonghoon Lee, Gary Geunbae Lee, and Minhwa Chung. Pronunciation variants prediction method to detect mispronunciations by Korean learners of English. ACM Transactions on Asian Language Information Processing, 13(4):16:1–16:??, December 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Liu:2014:DTL

[233] Lemao Liu, Tiejun Zhao, Taro Watanabe, Hailong Cao, and Conghui Zhu. Discriminative training for log-linear based SMT: Global or local methods. ACM Transactions on Asian Language Information Processing, 13(4):17:1–17:??, December 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Zhuang:2014:EPP

[234] Yi Zhuang, Qing Li, Dickson K. W. Chiu, Zhiang Wu, and Haiyang Hu. Efficient personalized probabilistic retrieval of Chinese calligraphic manuscript images in mobile cloud environment. ACM Transactions on Asian Language Information Processing, 13(4):18:1–18:??, December 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).